x=y-8 and -x-y=0 you suppose to be solving system using substitution
1
Expert's answer
2012-01-31T08:44:01-0500
Substitute x which is given by the first equation( x=y-8) into the second equation:
-(y-8)-y=0
now open the braces . As long as "-" is before the braces we should change the sign of terms inside the braces when opening them:
-y+8-y=0
-2y+8=0
2y=8
y=4
now find x from the first equation by substituting y=4:
x=4-8=-4
Thus the answer is x=-4, y=4